OpenStudy (anonymous):

1/3 * e^3x-1/3*e^-3X

OpenStudy (earthcitizen):

could you re-write with parenthesis ?

OpenStudy (anonymous):

[(1/3)e^3x]-[(1/3)e^-3x] x=2

OpenStudy (earthcitizen):

\[1/3(e ^{3(2)}-e^{(-3(2)})= 134.5\]

OpenStudy (earthcitizen):

or do the exponentials equate to 2 ?

OpenStudy (anonymous):

no I think you set it up right. thats how i set it up i was just confused on where to go from there but thank you !

OpenStudy (earthcitizen):

yh, no p

OpenStudy (anonymous):

I think I should give up to guessing around: [(1/3)e^3x] - [(1/3)e^-3x] x = 2 (1/3) (e^3x - x/e^3x = 2 (e^3x)^2 - x - 6e^3x = 0